One number is 6 more than another. If the sum of the smaller number and 3 times the larger number is 34, find the two numbers.
Answer by funmath(2933)

One number is 6 more than another. If the sum of the smaller number and 3 times the larger number is 34, find the two numbers.
Let the smaller number be: x
Then the larger number is 6 more than x: x+6
Sum means add
3 times means 3*
is means =
smaller number + 3*larger number = 34
(x)+3(x+6)=34
x+3x+18=34
4x+18=34
4x+18-18=34-18
4x=16
4x/4=16/4
x=4
The smaller number: x=4
The larger number: x+6=4+6=10
:
Check:
Does smaller number + 3*larger number = 34?
(4)+3(10)=34?
4+30=34?
34=34!
We're right!
Happy Calculating!!!